OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 D. -12

OpenStudy (nurali):

using the slope equation (y2-y1)/(x2-x1) = slope (-3-6)/(10-x) = -3/2 -9/10-x=-3/2 Cross Multiply -9*2=-3(10-x) -18=-30+3x -18+30=3x 3x=12 Divide by 3 both side x=4

OpenStudy (anonymous):

thank you so much! :)